Carolina starts on a 58 mile trip on her moped at 20 mph. After a while the motor kicks out and she pedals the remainder of the trip at 12 mph. The entire trip took 3 1/2 hours. How far did Carolina travel when the motor on the moped quit running?
Answer by mananth(16946)

riding time =x
pedalling time = y
x+y = 7/2
2x+2y=7..............1
20x+12y=58...........2
...

multiply (1)by -6
Multiply (2) by 1
-12 x + -12 y = -42
20 x + 12 y = 58

Add the two equations

8x=16

/8

x=2

plug value of x in (1)

2x+2y=7
4+2y=7
2y=7-4
2y=3
y=1.5
X=2
Distance = 2* 20 = 40 miles
it kicked out